Suppose that in 2016, the CPI for energy rose from 183.4 to 193.3 while the CPI for all items rose from 236.5 to 241.4. As a result the inflation rate for energy is lower than the overall inflation rate in 2016.

Select one:

True

False

For any given year, the CPI is the price of the basket of goods and services in the given year divided by the price of the basket in the base year, then multiplied by 100.

Select one:

True

False

Because the CPI is based on a fixed basket of goods, the introduction of new goods and services in the economy causes the CPI to overestimate the cost of living.

Select one:

True

False

The consumer price index (CPI) will be more influenced by a 10 percent increase in the price of food and beverages than a 10 percent increase in the price of housing

Select one:

True

False
Answer #1

- false
Inflation rate for energy = (193.3-183.4)*100/183.4= 5.4
Inflation rate at overall level = (241.4-236.5)*100/236.5= 2.07

- true

- true

- false

Housing has higher weightage so cpi will be affected more by increase in housing.